OpenStudy (anonymous):

factored form of 10a + a^2=0 explain.

OpenStudy (mathmath333):

example factored form of \(12=2\times 2\times 3\) like this \(10a^2+a=0------\text{take 'a' common}\\ \Large a(10a+1)=0\)

OpenStudy (mathmath333):

sry this one \(10a+a^2=0\\ \Large a(10+a)=0\)

OpenStudy (anonymous):

thank you :D

OpenStudy (mathmath333):

:D

OpenStudy (anonymous):

here ill tag you in another one

OpenStudy (mathmath333):

:C

OpenStudy (mathmath333):

ok

OpenStudy (anonymous):

sorry i didnt know if you wanted new medals or not

OpenStudy (anonymous):

so thats why im going to retag you

OpenStudy (mathmath333):

lol

OpenStudy (mathmath333):

u the 'right to tag'

OpenStudy (mathmath333):

i mean u have the right to tag

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

typo